Another busy couple of days at the New Forest Observatory over the bank holiday weekend working on the mini-WASP astrophotography array.

First job was upgrading the automatic computer-controlled dome rotator on the Pulsar 2.2m telescope observatory. Our first prototype used two 0.5Nm torque stepper motors. Whilst this proved that the idea worked, they didn't quite have the speed or torque required. Greg purchased some 2.2Nm Nema 23 hybrid stepper motors designed for CNC use, and looked very beefy and up to the job.

I wired up the new motors to the Stepperbee controller board and proved that the motors worked in the study. We then did all the metal-work required to mount the new motors in place of the old motors.

After some problems with earth-loops and truculent computers we got the new motors working properly with the observatory computers. My software runs on the the observatory computer and communicates with TheSKY to fetch the telescope position and sends the required signals to the motor controllers. A solid state compass provides closed loop positional feedback.

The motors easily slewed the dome at 3 or 4 degrees per second. We were able to run it even faster, but it is a bit terrifying at high speed - we don't want to lob the dome into the next garden! Here is a video of the system in action. I was very pleased with the speed!

This activity managed to use up most of Sunday afternoon. The skies looked very promising, so we set out to try some imaging. Whilst Greg concentrated on setup work in the old South Dome, I ran the mini-WASP array in the North dome, working my way through various problems, but was able to grab a few dozen frames of some nice targets with both cameras operating in parallel. Bed about 3am.

After a late start on Monday we tidied up a few problems and set to work with some stolen sheets and shirts to try and get a decent flat field. The best arrangement consisted of a couple of layers of white shirt attached to the Sky 90s with elastic bands pointing at the open telescope dome slit. A large thick white sheet was draped over the entire dome slit - providing a double diffused light source which provided a decent flat. Calibration with these flats and some bias frames removed all of the vignetting on the images from the previous night.

Next job is to look at the gigabyte or so of FITs files and see if there is a decent image lurking in them!